A coaxial cable consists of a solid inner conductor of radius a =0.2 m carrying total current of 5A, surrounded by a concentric cylindrical tube of inner radius b=0.4 m and outer radius c=0.6m carrying current of 5A. The conductors carry equal and opposite currents I distributed uniformly across their cross-sections. Determine the magnetic field at a distance r=0.5m from the axis.  (10p)



1
Expert's answer
2021-06-22T10:10:42-0400

Gives

a=0.2 m

Current=5A

b=0.4m

c=0.6m

r=0.5m



We know that

Ampear law in closed loop

"\oint B .dl=\mu_0I_{enclosed}"

"I_{en}=I_0"

"\oint B.dl=\mu_0(I-I_0)"

"\oint dl=2\pi r"

Area "\pi(c^2-b^2)" Flow current="I"

Area 1m2flow current"=\frac{I}{\pi(c^2-b^2)}"

Area "\pi(r^2-b^2)" Flow current"(I_0)" "=\frac{I}{\pi(c^2-b^2)}\times\pi(r^2-b^2)"


"I_0=\frac{\pi (r^2-b^2)}{\pi(c^2-b^2)}I"

"I_0=\frac{(r^2-b^2)}{(c^2-b^2)}I"

"B=\frac{\mu_0I}{2\pi r}\times(1-\frac{r^2-b^2}{c^2-b^2})"

Put value

Where

"\mu_0=4\pi\times10^{-7} Wb/A -m"

"B=\frac{4\times3.14\times10^{-7}\times5}{2\times3.14\times 0.5}\times(1-\frac{0.5^2-0.4^2}{0.6^2-0.4^2})"


"B=2\times10^{-6}\times0.55=1.1\times10^{-6}T"